Journalist Report – March 1th

Sol 17 – Saving Private LOAC

“Somewhere above, beyond, far off, was the sun.”

– – Chapter 17 of The Martian Chronicles by Ray Bradbury

I think Mars wants us off its turf… we awoke once again to the absurdity of snow in a warm, red desert. Our EVA to Candor Chasma was effectively cancelled, leaving Quentin and Alice, responsible for the photogrammetry experiment, with somewhat of a sinking feeling. Without the perspective of this EVA, it was a bit harder to get ourselves started, but it was quickly decided that we would get ahead and do our morning tasks for Sol 18 in advance. Corentin and Jérémy therefore started an EchoFinder session while Alice took a cognitive test in the Hab.

As the morning advanced and the temperatures rose, the thin layer of snow quickly melted, allowing us to conduct an emergency EVA to save the LOAC, an instrument prone to damage due to humidity, and change the other instruments’ batteries. In less than 30 minutes we were suited up: Alice, Adrien, and I would be guided by Quentin as HabCom, who could also give us technical advice from the Hab. While we managed to change the batteries, retrieve the data and reset the instruments, the field mill could not be rebooted, and the weather station was soaked: the screen stayed blank, no matter how hard I tried… We took the latter back inside the Hab to be inspected, as well as the LOAC.

Over lunch, we discussed our childhood reading habits, a conversation well suited to a rainy day. We were glad to have brought the LOAC back in before it started raining and snowing again by mid-afternoon: Alexandre took good care of him, letting him dry out and delicately brushing off any Martian dust that could have messed with the electronics. Restless as ever, I could not help but keep looking out the window to witness what I could not act upon. While few things are as frustrating to me as not being able to act on a problem, Alice is affected by the weather more than most, as she is in charge of both the photogrammetry and geology EVAs, and has to constantly adapt an already very complex schedule.

On a more positive note, Quentin has written some code to plot the different environmental data as functions of time, to correlate them with our activities within the station. He agreed to explain the graphs included in today’s batch of photos:

“ This is the evolution of 3 environmental parameters (temperature, humidity and light) as a function of time in the five rooms of the MDRS. By linking this information with physiological parameters, researchers can deduce how stressful a particular environment is.

But they can also simply reflect the activity of the Crew members during the day! In rectangle 1, we can see the evolution of the temperature in the GreenHab, which is heated during night. The temperature decreases, until the lower limit is attained, the heater is then switched on, and the temperature increases. And every half hour or so, this cycle is repeated!

In the rectangle 2, we can see the brightness increasing in the Lower Deck (yellow curve): somebody has switched on the light. But even more interesting: the humidity is increasing as well. Indeed, it’s time for our daily workout in the Lower Deck! ”

On this very Earth-like rainy day, I think it’s important to mention our very own “Mission Support” back on Earth. Every day, they work hard so that these very reports may reach their destination. Message from Crew 275 to Crew 293: Thank you for being our link to our home planet, and for occasionally sending in soccer and rugby scores…

EVA Report – March 1th

Crew 275 EVA Report 01Mar2023

EVA # 14

Author: Adrien Tison, EVA Leader and Crew Botanist

Purpose of EVA: The purpose of the EVA is to perform the photogrammetry experiments with the 3D map

First, as it is on the way to Candor Chasma, we will change the batteries and collect the data of MegaAres, LOAC, Purple Air, METEO and the Field Mill.

We will then go to Candor Chasma where the two subjects of the experiment will try to find some specific checkpoints, while the operator of the experiment onsite will take time measurements and draw their trajectory on a map. In addition, the two subjects will be equipped with a step counter.

Due to the bad weather conditions (snow), the purpose of the EVA was modified, in agreement with Mission Support, so that very urgent operations could be conducted with our atmospheric instruments.

New purpose of the EVA: Change the batteries and retrieve data from our atmospheric instruments (LOAC, MegaAres, the Weather station, Purple air, and the Field Mill), repatriate LOAC to the Hab.

Start time: 11:26

End time: 12:14


11:26 – Airlock closed.

11:34 – Rovers Spirit and Curiosity unplugged.

11:35 – Rovers on their way to the location.

11:37 – Rovers parked near the atmospheric instruments.

11:47 – Data retrieved, and battery changed for Mega-Ares.

11:49 – LOAC has been put in a box to take it back to the Hab.

11:54 – The zip-lock in which was the Weather station device was open, the LED stopped blinking, we then decided to take it back to the Hab too.

11:54 – Data retrieved, and battery changed for Purple Air.

11:59 – Data has been successfully retrieved from the Field Mill. However, there were some issues while trying to change the battery. Measurements didn’t start as expected and the device switched off.

12:04 – It started snowing so we couldn’t stay enough time to solve the issue, we then decided to go back to the rovers.

12:07 – Rovers loaded and ready to go.

12:11 – Arrival in front of the Hab.

12:12 – Loading materials into the Airlock.

12:14 – Airlock closed.

12:19 – End of the EVA.

Destination: Between the Hab and Marble Ritual near the road.

Coordinates: N4251000 E518500

Participants: EVA Leader: Adrien Tison (Crew Botanist), Alice Chapiron (Crew Scientist), and Marie Delaroche (Crew Journalist)

HAB COM: Quentin Royer (Crew Engineer)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual

Mode of travel: Driving using Spirit and Curiosity and walking

GreenHab Report – March 1th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 69.1 °F

Average humidity: 27%

Hours of supplemental light: 0 hour

Daily water usage for crops: 8.45 gallons

Daily water usage for research and/or other purposes: 2.11 gallons

Water in the blue tank: 138.17 gallons

Time(s) of watering for crops: 9:15

Changes to crops: N/A

Narrative: Today, as it was neither sunny nor hot, plants and crops didn’t suffer from heat and looked better. I still watered them well and took care of everything.

I noticed that the aquaponics system loses a certain amount of water in the long term. Indeed, mine is quite small so if you consider water loss due to the cycle, water volume decreases over time. I then added 2.11 gallons of water to the fish tank to ensure their safety and well-being for the following days.

However, the system is doing well, and you can easily notice how well the lettuce is growing!

Harvest: N/A

Support/supplies needed: N/A

Operations Report – March 1th

Crew 275 Operations Report 01Mar2023
SOL: 17

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• GreenHab Carbon Monoxide alarm

Notes on non-nominal systems: None

Spirit rover used: Yes
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: 91%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.7
• Beginning charge: 93%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.
• WATER USE: 18.02
• Water: 428.47 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 10.56 gallons
• Heater: On
• Supplemental light: No
• Harvest: No

Summary of ScienceDome operations: The Crew Scientist conducted experiments on rock samples.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: None.

Sol Summary Report – March 1th

Crew 275 Sol Summary Report 01Mar2023

Sol: 17

Summary Title: Saving Private LOAC

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

The weather surprised us again today and we had to cancel our exploration EVA to Candor Chasma, initially planned in the morning. When it stopped snowing, Mission Support agreed to let us put on our spacesuits to quickly change the batteries of our atmospheric instruments and retrieve the LOAC device, as it is very sensitive to precipitation. With the rainy afternoon that we had, we were indeed happy that the LOAC was safe within the Hab!

All these changes in the schedule give additional work to our Crew Scientist, who is responsible for planning our scientific activities. At least, we all hope that the ground will soon be dry so that we can go back on EVA, especially during the visit of the BBC on Friday (Sol 19).

Today, we have finished equipping everyone with the tracking system, so that we are aware of our position at anytime of the day and correlate this with other information. We have also continued the human factors protocols, as well as the tests of the ultrasound device based on virtual reality and artificial intelligence.

Look Ahead Plan:

We would like to reschedule today’s EVA at Candor Chasma. The plan would be to have a first EVA on Friday (Sol 19) and another one on Saturday (Sol 20). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render, respectively. We would like to take the opportunity of these EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Rainy, low 0°C/32°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: The objectives of EVA #14 were changed due to the bad weather conditions. This EVA will have to be rescheduled.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site